OpenStudy (anonymous):

Crusty the clown is pushing a 4.00x10^3 N piano with a force of 510 N across a rough floor which causes a friction of 405 N. What speed will the piano have after 5.0s if it starts from rest?

OpenStudy (anonymous):

F(weight) = 4000 N F(applied) = 510 N F(friction) = -405 N t = 5 s g = 9.81 m/s^2 F(weight) = mg m = 4000/9.81 = 407.75 kg The net force acting on the piano is: F(net) = F(applied) + F(friction) = 510 - 405 = 105 N F = ma a = 105/407.75 = 0.258 m/s^2 v = at v = (0.258)(5) = 1.29 m/s Answer: 1.29 m/s